# The dynamical and thermodynamic effects of turbulence for the cosmic baryonic fluid

### v1

@@
Abstract
Both simulations and observations indicate that the so-called missing baryons reside in the intergalactic medium (IGM) known as the warm-hot intergalactic medium (WHIM). In this article, we demonstrate that turbulence in the cosmic baryonic fluid is crucial for correctly understanding both the spatial distribution and the physical origins of the missing baryons in the universe. First, we find that dynamical effects cause the gas to be detained in low-density and intermediate-density regions, resulting in high baryon fractions, while prevent the inflow of the gas in high-density regions, leading to low baryon fractions. Second, turbulent energy is converted into thermal energy, and the injection and dissipation of turbulent energy have essentially reached a balance from $z=1$ to 00. This indicates that the cosmic fluid is in a state of fully-developed turbulence within this redshift range. Due to turbulent heating, as redshift decreases, an increasing amount of warm gas is heated and transitions into the WHIM, and some even into hot gas.
@@ 1 Introduction
According to the standard cosmological model, approximately 5% of the total energy density of the universe consists of baryonic matter, which includes protons and neutrons. However, observations of stars, galaxies, and the surrounding gas have shown that about $30\%-40\%$ of the predicted baryons are missing (Fukugita et al., 1998; Bregman, 2007; Fang, 2018). Scientists speculate that these missing baryons may reside in a low-density high-temperature plasma, known as the WHIM (Cen & Ostriker, 1999; Davé et al., 2001; Fang et al., 2002, 2006; Branchini et al., 2009).
@@
In recent years, with the advancement of new astronomical observation techniques and methods, scientists have been able to detect some of the previously undetectable baryons through X-ray and radio facilities (Nicastro et al., 2018; Nevalainen et al., 2019; Macquart et al., 2020). These studies support the idea that these baryons in the universe are not truly missing but are present in the IGM within the cosmic web structures. Despite significant progress, the precise locations and physical origins of the missing baryons remain elusive (Driver, 2021).
Based on our previous series of works (Kim et al., 2005; He et al., 2006; Yang et al., 2020, 2022; Wang & He, 2024a, b), in this work, we demonstrate that turbulence in the cosmic baryonic fluid is crucial for correctly understanding both the spatial distribution and the physical origins of the missing baryons in the universe. The consensus reached is that turbulence can occur at any scale in the cosmic fluid, which has been extensively studied both theoretically (Norman & Bryan, 1999; Brüggen et al., 2005; Cassano & Brunetti, 2005; Dolag et al., 2005; Subramanian et al., 2006; Roediger & Brüggen, 2007; Ryu et al., 2008; Lau et al., 2009; Zhu et al., 2010; Gaspari et al., 2011; Evoli & Ferrara, 2011; Vazza et al., 2012; Zhu et al., 2013; Banerjee & Sharma, 2014; Miniati, 2014; Porter et al., 2015; Iapichino et al., 2017; Angelinelli et al., 2020) and observationally (Vogt & Enßlin, 2003; Schuecker et al., 2004; Murgia et al., 2004; Enßlin & Vogt, 2006; Bonafede et al., 2010; Vacca et al., 2010; Churazov et al., 2012; Zhuravleva et al., 2014; Walker et al., 2015; Khatri & Gaspari, 2016; Shi & Zhang, 2019). Through dynamical and thermodynamic investigations, we provide a detailed analysis of how turbulence influences the distribution of baryonic matter in cosmic space and offer a theoretical explanation for the missing baryon problem.
The paper is organized as follows. In Section 2, we briefly introduce our methods used in this work. In Section 3, we present the results of our work. In Section 4, we present the conclusions and discussions of the paper.
@@ 2 Methods
We divide the space into the following regions, according to the dark matter density $\Delta_{\rm dm}({\bf x})=\rho_{\rm dm}({\bf x})/{\bar{\rho}}_{\rm dm}$, as (1) low-density, with $\Delta_{\rm dm}({\bf x})<0.2-0.5$, depending on redshifts, (2) intermediate-density, with $2000.2-0.5<\Delta_{\rm dm}({\bf x})<$ 200, (3) high-density, with $200<\Delta_{\rm dm}({\bf x})<10^{3}$, and (4) extremely high-density, with $\Delta_{\rm dm}({\bf x})>10^{3}$, which roughly correspond to (1) voids and under-dense regions, (2) filaments, sheets and outskirts of clusters on the cosmic web, (3) virialized structures such as galaxy clusters, and (4) galaxies in clusters. This classification scheme is roughly consistent with that of Vazza et al. (2009).
Similar to Bregman (2007), we also define 4 components of baryonic matter, as (1) hot gas, with $T>10^{7}$, (2) WHIM, with $10^{5}$$<T<10^{7}$, (3) warm gas, with $T<10^{5}$, and (4) cold component, with star, AGN (black hole) and (possible) others.
@@ 2.1 Divergence of velocity field and dynamical effects
in which $a$ is the scale factor, and the thermal pressure $P$ can be computed using equation (10). $Q_{\rm turb}$, $Q_{\rm th}$, $Q_{\rm grav}$ and $Q_{\rm exp}$ indicate the dynamical effects of turbulence, thermal, gravitational collapse, and cosmic expansion, respectively, and $Q_{\rm tot}=Q_{\rm turb}+Q_{\rm th}+Q_{\rm grav}+Q_{\rm exp}$ represents their total combined effects. In the above expression, $\bm{\omega}\equiv\nabla\times{\bf u}$ represents the vorticity, and $S_{ij}\equiv(1/2)({\partial_{i}}{u_{j}}+{\partial_{j}}{u_{i}})$ represents the strain rate of the velocity field for the cosmic fluid.
@@ 2.2 Mass fraction and baryon fraction
For the simulation data, we define the total mass fraction within the simulation volume as a function of redshift as
$f_{\rm x}(z)=\frac{M_{\rm x}(z)}{{M}_{\rm total}},$ (3)
where the subscript ‘${\rm x}$’ refers to the 4 baryonic components, $M_{\rm x}(z)$ denotes the total mass of the ${\rm x}$-component at redshift $z$, and $M_{\rm total}$ refers to the total baryonic mass within the simulation box.
We specifically designate a sequence of dark matter densities as follows:
@@
For the spatial points within the $i$-th dark matter density interval at redshift $z$, we count the masses of the 4 baryonic components defined earlier, as $m_{\rm x}(\Delta_{i},z)$, and we define the cumulative mass fraction of the 4 baryonic components that depend on dark matter density as
[…]
Note that the summation in the above expression is performed in descending order from $N$ to $i$, that is, from high-density to low-density regions. The total cumulative mass fraction is defined by summation over the 4 baryonic components as
@@
We also need to define baryon fraction as a function of spatial positions and redshift for our work, as
@@
2.3 Turbulent and thermal energy of the baryonic field
TNG provides the internal energy per unit mass of baryonic gas, $e_{\rm u}({\bf x})$. Hence we have the thermal energy density,
where $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm p}$ the mass of the proton. $\mu({\bf x})$, $\rho_{\rm b}({\bf x})$ and $T({\bf x})$, dependent on the spatial locations, are the mean molecular weight, mass density, and temperature of the baryonic fluid, respectively. Regarding the cosmic fluid as the monoatomic ideal gas, its thermal pressure is
The turbulent energy density is
in which ${\bf u}_{\rm turb}({\bf x})$ is the three-dimensional velocity of turbulent flow. In general, a velocity field of the fluid ${\bf u}$ is a superposition of the turbulent velocity and the bulk velocity, i.e. ${\bf u}={\bf u}_{\rm turb}+{\bf u}_{\rm bulk}$. As in Wang & He (2024a, b), we also use the iterative multi-scale filtering approach developed by Vazza et al. (2012) to extract turbulent motions from the velocity field of the cosmic fluid.
We also define the ratio of turbulent energy to thermal energy as
For the random velocity field of the cosmic baryonic fluid ${\bf u}_{\rm turb}({\bf x})$ with the zero mean value, its isotropic continuous wavelet transform (CWT) $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$ is obtained by convolution with the wavelet function $\Psi$ as
$\tilde{{\bf u}}_{\rm turb}(w,{\bf x})=\int{\bf u}_{\rm turb}(\bm{\tau})\Psi(w,{\bf x}-\bm{\tau}){\rm d}^{3}\bm{\tau}.$ (13)
Throughout this work, we use the so-called 3D isotropic cosine-weighted Gaussian-derived wavelet (CW-GDW), which can achieve good localization in both spatial and frequency space simultaneously (Wang & He, 2022, 2024c, 2024a). With $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$, and referring to equation (11), we define the local wavelet power spectrum (WPS) for the turbulent velocity $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$ as
Note that here we do not use $\rho_{\rm b}({\bf x})$ directly, but instead use $\Delta_{\rm b}({\bf x})=\rho_{\rm b}({\bf x})/\bar{\rho}_{\rm b}$ for the calculations, where $\bar{\rho}_{\rm b}$ is the background baryonic density.
Then we define the environment-dependent WPS as
where the environment is specified with the density contrast $\delta$ and the mean $_{\delta({\bf x})=\delta}$ is performed over all the spatial points where the condition ${\delta({\bf x})=\delta}$ is satisfied.
For details of the CWT techniques that we have developed, please refer to Wang & He (2021); Wang et al. (2022); Wang & He (2022, 2023, 2024c).
Figure 1: Mean divergence $\overline{d({\bf x})}$ of the baryonic velocity field as a function of dark matter density. For a spatial location ${\bf x}$, there exists the correspondence between the divergence $d({\bf x})=\nabla\cdot{\bf u}({\bf x})$ of baryonic velocity field ${\bf u}$ and the dark matter density $\Delta_{\rm dm}({\bf x})=\rho_{\rm dm}({\bf x})/\bar{\rho}_{\rm dm}$. For a density interval centered at $\Delta_{\rm dm}$, we count all the points within this interval, and in this way, we can calculate the mean value and the standard deviation of $d({\bf x})$. The color band indicates the standard deviations for the values above and below the mean, which are calculated separately. For clarity, only the deviations for $z=0$ are displayed; the results for other redshifts are similar to those of $z=0$. Throughout the article, we use an overbar to indicate a statistically averaged quantity.
Figure 2: Mean dynamical effects as a function of dark matter density. The mean dynamical effects of turbulence and thermal motion in the cosmic fluid, as well as the gravitational effect, the cosmic expansion effect, and their combined effects, are represented by $Q_{\rm turb}$, $Q_{\rm th}$, $Q_{\rm grav}$, $Q_{\rm exp}$ and $Q_{\rm tot}$, respectively. Panels (a) to (f) are for $z=4$ to 00, respectively. The lines represent the mean values of $Q({\bf x})$, and the color bands indicate the standard deviations for the values above and below the mean, which are calculated separately.
Figure 3: Mass fraction as a function of redshift. The mass fractions of the 4 baryonic components evolve with redshift.

### v3

# The dynamical and thermodynamic effects of turbulence on the cosmic baryonic fluid

@@
11institutetext: College of Physics, Jilin University, Changchun 130012, China 22institutetext: Center for High Energy Physics, Peking University, Beijing 100871, China
The dynamical and thermodynamic effects of turbulence on the cosmic baryonic fluid
Yun Wang (王云)    Minxing Li (李敏行)    Ping He (何平)
Both simulations and observations indicate that the so-called missing baryons reside in the intergalactic medium known as the warm-hot intergalactic medium (WHIM). In this study we employed the IllustrisTNG50-1 simulation to demonstrate that knowledge of the turbulence in the cosmic baryonic fluid is crucial for correctly understanding both the spatial distribution and the physical origins of the missing baryons in the Universe. First, we find that dynamical effects cause the gas to be detained in low-density and intermediate-density regions, resulting in high baryon fractions, and prevent the convergence of the gas in high-density regions, leading to low baryon fractions. Second, turbulent energy is converted into thermal energy, and the injection and dissipation of turbulent energy have essentially reached a balance from $z=1$ to 0. This indicates that the cosmic fluid is in a steady state within this redshift range. Due to turbulent heating, as the redshift decreases, an increasing amount of warm gas is heated and converted into the WHIM, and some even into hot gas. We find that, compared with turbulence in the cosmic fluid, shocks are unimportant in intermediate-density regions and even negligible in high-density regions, both dynamically and thermodynamically. This finding accounts for the origin of the WHIM in terms of both dynamics and thermodynamics, calls into question the traditional view of shock-heating, and highlights the importance of turbulence in shaping the large-scale structure of the Universe, particularly in the evolution of galaxies and galaxy clusters. In addition to TNG50-1, we validated our key findings with TNG50-2, TNG100-1, WIGEON, and EAGLE simulations, demonstrating that the spatial resolution, box size, and sub-grid-physics variations do not affect our main conclusions.
Key Words.:
turbulence – methods: numerical – galaxies: clusters: intracluster medium – large-scale structure of Universe
@@ 1 Introduction
According to the standard cosmological model, approximately 5% of the total energy density of the Universe consists of baryonic matter, which includes protons and neutrons. However, observations of stars, galaxies, and the surrounding gas have shown that about $30\%-40\%$ of the predicted baryons are missing (Fukugita et al., 1998; Bregman, 2007; Fang, 2018). Scientists speculate that these missing baryons may reside in a low-density high-temperature plasma known as the warm-hot intergalactic medium (WHIM; Cen and Ostriker 1999; Davé et al. 2001; Fang et al. 2002, 2006; Branchini et al. 2009). In recent years, with the advancement of new astronomical observation techniques and methods, scientists have been able to detect some of the previously undetectable baryons through X-ray and radio facilities (Nicastro et al., 2018; Nevalainen et al., 2019; Macquart et al., 2020; Migkas et al., 2025). Through the analysis of dispersion measurements from fast radio bursts, Connor et al. (2025) determined that the majority of the missing baryons exist as diffuse ionized gas in the intergalactic medium (IGM). These studies support the idea that these baryons are not truly missing but are present in the IGM within the cosmic web structures. Despite significant progress, the precise locations and physical origins of the missing baryons remain elusive (Driver, 2021).
For a long time, shocks created by structure formation or accretion have been considered important thermal sources responsible for heating the IGM and the intracluster medium (ICM; Vazza et al. 2009a, 2010, 2011, 2013; Shull et al. 2012; Miniati 2014, 2015). Kim et al. (2005) and Zhang et al. (2006) show that the cosmic baryonic fluid in the nonlinear regime behaves like Burgers turbulence, i.e., it is characterized by shocks, which plays a significant role in converting kinetic energy into thermal energy, thereby heating the gas. Additionally, shocks have also been assumed to heat the warm gas and turn it into the WHIM in intermediate-density regions (Cen and Ostriker, 1999; Davé et al., 1999, 2001; Smith et al., 2011; Schmidt et al., 2016). However, our latest studies, based on IllustrisTNG50 cosmological simulation data, indicate that such a perspective needs to be revised. In this paper we demonstrate that turbulence is a more powerful heating mechanism than shocks.
Turbulent motions in the cosmic baryonic fluid within large-scale structures have attracted growing attention in cosmological studies in recent decades. A well-established consensus postulates that turbulence can arise across all scales in the cosmic fluid; this has been extensively studied both theoretically (Norman and Bryan, 1999; Brüggen et al., 2005; Cassano and Brunetti, 2005; Dolag et al., 2005; He et al., 2006; Subramanian et al., 2006; Roediger and Brüggen, 2007; Ryu et al., 2008; Lau et al., 2009; Zhu et al., 2010; Gaspari et al., 2011; Evoli and Ferrara, 2011; Vazza et al., 2012; Zhu et al., 2013; Miniati, 2014; Porter et al., 2015; Iapichino et al., 2017; Angelinelli et al., 2020) and observationally (Vogt and Enßlin, 2003; Schuecker et al., 2004; Murgia et al., 2004; Enßlin and Vogt, 2006; Bonafede et al., 2010; Vacca et al., 2010; Churazov et al., 2012; Zhuravleva et al., 2014; Walker et al., 2015; Khatri and Gaspari, 2016; Shi and Zhang, 2019). Studies of turbulence in the cosmic baryonic fluid carry profound implications for both astrophysics and cosmology, particularly in terms of the following aspects:
(1) Turbulence in baryonic fluid significantly influences the formation and evolution of cosmic large-scale structures (Zhu et al., 2010). Subsonic turbulence transfers and dissipates kinetic energy across different scales, while supersonic turbulence compresses gas to form shocks, thereby affecting the temperature, density distribution, and gravitational collapse processes of the gas (Chen et al., 2023; Zhuravleva et al., 2014). Research on the turbulent heating of baryonic fluid can more accurately simulate and predict the formation of cosmic structures, providing crucial evidence for refining cosmological models (Bauer and Springel, 2012; Springel, 2010).
(2) Turbulence in baryonic fluid is closely related to the formation and evolution of galaxies. It can affect gas cooling, star formation, and the nature of the interstellar medium (ISM) in galaxies (Howes, 2010). For instance, the thermal effects of turbulence can suppress gas cooling, thereby influencing the rate of star formation (Zubovas et al., 2024). Additionally, turbulence can drive the mixing and diffusion of gases, affecting the chemical composition and dynamical state of the ISM (Sharda et al., 2024). Turbulence heating has a significant impact on the distribution and motion of baryonic matter, particularly on small scales, and its effects are comparable to those of supernova (SN) and active galactic nucleus (AGN) feedback processes (Yang et al., 2020, 2022). Studying baryonic fluid turbulence helps deepen our understanding of the mechanisms of galaxy formation and their evolutionary history.
(3) Turbulence in baryonic fluid has a notable impact on the thermodynamic properties of galaxy clusters (Wang et al., 2023). The kinetic and thermodynamic effects of turbulence can influence the distribution of density, temperature, and entropy in galaxy clusters, as well as X-ray radiation (Banerjee and Sharma, 2014). By studying baryonic fluid turbulence, we can better explain observed phenomena such as temperature profiles and entropy distributions in galaxy clusters, thereby revealing their formation and evolutionary processes (Kay and Pratt, 2022).
(4) Turbulence in baryonic fluid is closely related to the acceleration and propagation of cosmic rays (Hanasz et al., 2021). Turbulent kinetic energy provides the energy for cosmic ray acceleration and also affects their propagation and distribution in the ISM and galaxy clusters (Armillotta et al., 2021). Research on the heating mechanisms of baryonic fluid turbulence helps us understand the origin of cosmic rays and their propagation mechanisms, as well as their impact on high-energy astrophysical processes in the Universe (Lazarian et al., 2023).
In a previous study (Wang and He, 2024c), we employed IllustrisTNG simulation data to investigate the turbulent and thermal motions of the cosmic baryonic fluid. With continuous wavelet transform (CWT) techniques, we defined the pressure spectra, or density-weighted velocity power spectra, as well as the spectral ratios, for both turbulent and thermal motions. We find that the magnitude of the turbulent pressure spectrum grows slightly from $z=4$ to $2$ and increases significantly from $z=2$ to $1$ at large scales, suggesting progressive turbulence injection into the cosmic fluid, whereas from $z=1$ to 0, the spectrum remains nearly constant, indicating that turbulence may be balanced by energy transfer and dissipation. The magnitude of the turbulent pressure spectra also increases with environmental density, with the highest-density regions showing a turbulent pressure up to six times that of thermal pressure. We also explored the dynamical effects of turbulence and thermal motions, discovering that while thermal pressure works to prevent structure collapse, turbulent pressure almost counteracts this effect, calling into question the common belief that turbulent pressure helps prevent gas from overcooling. We also find that, in the inertial range, the energy spectrum’s exponent is steeper than both the Kolmogorov and Burgers exponents, indicating more efficient energy transfer compared to Kolmogorov or Burgers turbulence (Wang and He, 2024b).
In this work we demonstrate that turbulence in the cosmic baryonic fluid is crucial for correctly understanding both the spatial distribution and the physical origins of the missing baryons in the Universe. Through dynamical and thermodynamic investigations, we provide a detailed analysis of how turbulence influences the distribution of baryonic matter in cosmic space and offer a theoretical explanation for the missing baryon problem. The paper is organized as follows. In Sect. 2 we briefly introduce the methods used in this work. In Sect. 3 we present our results. In Sect. 4 we present the conclusions and discussions.
@@ 2 Methods
We divided the space into the following regions according to the dark matter density, $\Delta_{\rm dm}({\bf x})=\rho_{\rm dm}({\bf x})/{\bar{\rho}}_{\rm dm}$: (1) low-density, with $\Delta_{\rm dm}({\bf x})<0.2-0.5$, depending on redshifts, (2) intermediate-density, with $0.2-0.5<\Delta_{\rm dm}({\bf x})<200$, (3) high-density, with $200<\Delta_{\rm dm}({\bf x})<10^{3}$, and (4) extremely high-density, with $\Delta_{\rm dm}({\bf x})>10^{3}$. This roughly correspond to (1) voids and under-dense regions, (2) filaments, sheets, and outskirts of clusters in the cosmic web, (3) virialized structures such as galaxy clusters, and (4) galaxies in clusters, respectively. This classification scheme is roughly consistent with that of Vazza et al. (2009a). Similar to Bregman (2007), we also defined four components of baryonic matter: (1) hot gas, with $T>10^{7}$K, (2) the WHIM, with $10^{5}$$T<10^{7}K$, (3) warm gas, with $T<10^{5}$K, and (4) the cold component, with stars, AGNs (black holes), and (possibly) other objects.
@@ 2.1 Divergence of velocity field and dynamical effects
Here $a$ is the scale factor, and the thermal pressure ($P$) can be computed using Eq. (4). $Q_{\rm turb}$, $Q_{\rm th}$, $Q_{\rm grav}$, and $Q_{\rm exp}$ indicate the dynamical effects of turbulence, thermal, gravitational collapse, and cosmic expansion, respectively, and $Q_{\rm tot}=Q_{\rm turb}+Q_{\rm th}+Q_{\rm grav}+Q_{\rm exp}$ represents their total combined effects. In the above expression, $\boldsymbol{\omega}\equiv\nabla\times{\bf u}$ represents the vorticity, and $S_{ij}\equiv(1/2)({\partial_{i}}{u_{j}}+{\partial_{j}}{u_{i}})$ represents the strain rate of the velocity field for the cosmic fluid.
2.2 Turbulent and thermal energy of the baryonic field
IllustrisTNG provides the internal energy per unit mass of baryonic gas, $e_{\rm u}({\bf x})$. Hence, we have the thermal energy density,
where $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm p}$ the mass of the proton. $\mu({\bf x})$, $\rho_{\rm b}({\bf x}),$ and $T({\bf x})$, dependent on the spatial locations, are the mean molecular weight, mass density, and temperature of the baryonic fluid, respectively. Regarding the cosmic fluid as the monoatomic ideal gas, its thermal pressure is
The turbulent energy density is
in which ${\bf u}_{\rm turb}({\bf x})$ is the 3D velocity of turbulent flow. In general, a velocity field of the fluid $\bf u$} is a superposition of the turbulent velocity and the bulk velocity, i.e., ${\bf u}={\bf u}_{\rm turb}+{\bf u}_{\rm bulk}$. As in Wang and He (2024b, c), we also used the iterative multi-scale filtering approach developed by Vazza et al. (2012) to extract turbulent motions from the velocity field of the cosmic fluid.
We defined the ratio of turbulent energy to thermal energy as
For the random velocity field of the cosmic baryonic fluid ${\bf u}_{\rm turb}({\bf x})$ with the zero mean value, its isotropic CWT $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$ is obtained by convolution with the wavelet function $\Psi$ as
$\tilde{{\bf u}}_{\rm turb}(w,{\bf x})=\int{\bf u}_{\rm turb}(\boldsymbol{\tau})\Psi(w,{\bf x}-\boldsymbol{\tau}){\rm d}^{3}\boldsymbol{\tau},$ (7)
in which $\Psi(w,\mathbf{x})=w^{3/2}\Psi(w|{\bf x}|)$, with $w=0.3883k$ for the 3D isotropic cosine-weighted Gaussian-derived wavelet (CW-GDW). Throughout this work, we used the CW-GDW, which can achieve good localization in both spatial and frequency space simultaneously (Wang and He, 2022, 2024a, 2024b). With $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$, and referring to Eq. (5), we defined the local wavelet power spectrum (WPS) for the turbulent velocity $\tilde{{\bf u}}_{\rm turb}(w,{\bf x})$ as
Note that here we do not use $\rho_{\rm b}({\bf x})$ directly, but instead use $\Delta_{\rm b}({\bf x})=\rho_{\rm b}({\bf x})/\bar{\rho}_{\rm b}$ for the calculations, where $\bar{\rho}_{\rm b}$ is the background baryonic density. Then we defined the environment-dependent WPS as
where the environment is specified with the dark matter density contrast $(\delta$) and the mean $\left<...\right>_{\delta({\bf x})=\delta}$ is performed over all the spatial points where the condition $\delta({\bf x})=\delta$} is satisfied. For details of the CWT techniques that we have developed, please refer to Wang and He (2021, 2022, 2023, 2024a) and Wang et al. (2022).
2.3 Contributions of turbulence and shocks
There are many shock-finding algorithms in current literature (e.g., Pfrommer et al., 2006; Skillman et al., 2008; Vazza et al., 2009a, b; Beck et al., 2016; Schaal and Springel, 2015; Schaal et al., 2016; Valdarnini, 2019; Vallés-Pérez et al., 2021), which are basically designed based on Rankine-Hugoniot shock jump conditions. Generally, when the fluid velocity exceeds the local speed of sound, the fluid’s compression and expansion processes cannot be balanced through sound wave propagation, leading to a discontinuity in the fluid state, i.e., a shock wave. As usual, the supersonic motion is described by the Mach number $\mathcal{M}\equiv u/c_{s}>1$, where $u$ is the velocity of the fluid field and $c_{s}$ is the sound speed. In the current work, we used the shock finding algorithm of Schaal and Springel (2015) and Schaal et al. (2016) to identify shock zones. For a small dark matter density interval centered at $\Delta_{\rm dm}$, we counted a total of $n_{\rm g}$ grid points, and for every spatial position $\bf x$}, we assigned a $m({\bf x})$ number of $1$ or 0, according to whether it is marked as shock or turbulence. Then, the averaged dynamical effects contributed by shocks and turbulence could be derived as
$\overline{Q}_{\rm shock}(\Delta_{\rm dm})$ $=\frac{1}{n_{\rm g}}\sum^{n_{\rm g}}_{i=1}m({\bf x}_{i})Q_{\rm turb}({\bf x}_{i}),$
$\overline{Q}_{\rm turb}(\Delta_{\rm dm})$ $=\frac{1}{n_{\rm g}}\sum^{n_{\rm g}}_{i=1}(1-m({\bf x}_{i}))Q_{\rm turb}({\bf x}_{i}),$ (10)
in which $Q_{\rm turb}({\bf x})$ is computed with Eq. (2.1). The averaged thermal effects contributed by shocks and turbulence, i.e., the heating of the gas, can be derived as
$\overline{\varepsilon}_{\rm shock}(\Delta_{\rm dm})$ $=\frac{1}{n_{\rm g}}\sum^{n_{\rm g}}_{i=1}m({\bf x}_{i})\varepsilon({\bf x}_{i}),$
$\overline{\varepsilon}_{\rm turb}(\Delta_{\rm dm})$ $=\frac{1}{n_{\rm g}}\sum^{n_{\rm g}}_{i=1}(1-m({\bf x}_{i}))\varepsilon_{\rm turb}({\bf x}_{i}),$ (11)
in which $\varepsilon({\bf x})=\frac{1}{2}\rho_{\rm b}({\bf x}){\bf u}^{2}({\bf x})$ is computed with ${\bf u}={\bf u}_{\rm turb}+{\bf u}_{\rm bulk}$ and $\varepsilon_{\rm turb}({\bf x})$ is with Eq. (5).
Figure 1: Mean divergence, $\overline{d({\bf x})}$, of the baryonic velocity field as a function of dark matter density. For a spatial location $\bf x$}, there exists the correspondence between the divergence, $d({\bf x})=\nabla\cdot{\bf u}({\bf x}),$ of baryonic velocity field ($\bf u$}) and the dark matter density, $\Delta_{\rm dm}({\bf x})=\rho_{\rm dm}({\bf x})/\bar{\rho}_{\rm dm}$. For a density interval centered at $\Delta_{\rm dm}$, we counted all the points within this interval, which enabled us to calculate the mean value of $d({\bf x})$. Throughout the paper, we use an overbar to indicate a statistically averaged quantity.
@@ 2.4 Mass fraction and baryon fraction
For the simulation data, we defined the total mass fraction within the simulation volume as a function of redshift as
$f_{\rm x}(z)=\frac{M_{\rm x}(z)}{{M}_{\rm total}}.$ (12)
The subscript $\rm x$}” refers to the four baryonic components, $M_{\rm x}(z)$ denotes the total mass of the $\rm x$}-component at redshift $z$, and $M_{\rm total}$ refers to the total baryonic mass within the simulation box. We specifically designated a sequence of dark matter densities as follows:
@@
For the spatial points within the $i$-th dark matter density interval at redshift $z$, we counted the masses of the four baryonic components defined earlier, as $m_{\rm x}(\Delta_{i},z)$, and we defined the cumulative mass fraction of the four baryonic components that depend on dark matter density as
$F_{\rm x}(\Delta_{i},z)=\frac{1}{M_{\rm total}}\sum_{k=N}^{i}m_{\rm x}(\Delta_{k},z),\ \ 1\leq i\leq N.$ (15)
Note that the summation in the above expression is performed in descending order from $N$ to $i$, that is, from high-density to low-density regions. The total cumulative mass fraction is defined by summation over the four baryonic components as
@@
We also needed to define the baryon fraction as a function of spatial positions and redshift for our work:
@@
Figure 2: Mean dynamical effects as a function of dark matter density. The mean dynamical effects of turbulence and thermal motion in the cosmic fluid, as well as the gravitational effect, the cosmic expansion effect, and their combined effects, are represented by $Q_{\rm turb}$, $Q_{\rm th}$, $Q_{\rm grav}$, $Q_{\rm exp}$, and $Q_{\rm tot}$, respectively. Panels (a) to (f) are for $z=4$ to 0, respectively. We chose the units of velocity, $\nabla$, and time as ${\rm km}/s$, $h/{\rm Mpc}$, and $({\rm Mpc}/h/{\rm km})s$, respectively, which results in the vertical label shown in the figure.
Figure 3: Mean energy densities and energy ratios as a function of dark matter density. The dimensionless baryonic density $\Delta_{\rm b}\equiv\rho_{\rm b}({\bf x})/\bar{\rho_{\rm b}}$ is used to compute the energy density. Panels (a) and (b) show the mean energy densities of thermal and turbulent energy at different redshifts, respectively. Panel (c) shows the energy ratios of turbulence to thermal energy. The lines represent the mean values.
Figure 4: $z$-evolution of environment-dependent wavelet energy spectra of turbulence. Panels (a), (b), and (c) show the energy spectra of three intervals of dark matter density ($\Delta_{\rm dm}$), as marked. The energy spectra $E_{\rm turb}(k,\delta)$ are obtained from the env-WPS in Eq. (9) as $E_{\rm turb}(k,\delta)=k^{2}S_{\rm turb}(k,\delta)$. The relevant power-law fits for the scale ranges that are smaller and larger than the peak positions ($k_{\rm peak}$) are indicated.